OpenStudy (anonymous):

What is the value of the missing exponent? (Type only the numerical value of the missing exponent. Do not skip any spaces.) (-2x^3)?=-8x^9

OpenStudy (pokemon23):

Let me put my thinking cap on...

OpenStudy (anonymous):

What is the value of the missing exponent

OpenStudy (pokemon23):

I really don't know...? I hate word problems

OpenStudy (pokemon23):

Let me ask people from the math group

OpenStudy (pokemon23):

hold I will get you help

OpenStudy (anonymous):

Okay thank you very much

OpenStudy (pokemon23):

Ok Mr.Math is a genius in math

OpenStudy (mr.math):

Alright. We want to find a here \[(-2x^3)^{a}=-8x^9 \] \[(-2x^3)^a=-2^3(x^3)^3\] \[ (-2x^3)^a=(-2x^3)^3\]

OpenStudy (mr.math):

What do you think a should be here?

OpenStudy (mr.math):

So both sides of the equation are equal.

OpenStudy (pokemon23):

thanks Mr.Math

OpenStudy (mr.math):

You're very welcome!

OpenStudy (anonymous):

Thank you Mr.Math and Pokemon23 you've been very helpful.